Circular No. 7880

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1901+014 R. A. Remillard, Massachusetts Institute of Technology (MIT); and

D. A. Smith, University of Michigan, for the RXTE ASM Team at MIT and NASA Goddard Space Flight Center, report the discovery of a fast x-ray transient. The detection was initially thought to be a γ-ray burst

(cf. http://gcn.gsfc.nasa.gov/gcn/gcn3/1332.gcn3), with a peak flux

of 0.9 Crab (2–12 keV). The event began on Apr. 6.76268 UT, lasting > 2 min and < 3.15 hr (see http://xte.mit.edu/xtej1901+104). The ASM data show rapid evolution to a very hard spectrum during this outburst, but the IPN instruments failed to detect the source above 15 keV (K. Hurley, private communication). Furthermore, archival ASM data show a previous outburst at the same position on 1997 June 21.215, which was longer than 6 min and < 8 hr, with a peak flux of 0.4 Crab. The combined ASM observations yield a best position α = 19^{h}01^{m}36^{s}, δ = +1^{o}26′.4 (equinox J2000.0; estimated uncertainty 3′ in α and 2′ in δ). The error box contains one very faint, unidentified x-ray source from the Einstein Galactic Plane Survey: 2E 1859.1+0122 (α = 19^{h}01^{m}40^{s}.2, δ = +1^{o}27′13′′, equinox J2000.0; error radius 48′′; Hertz and Grindlay 1988, A.J. 96, 233). The time scales for these eruptions are reminiscent of the microquasar and blackhole binary V4641 Sgr, so optical and radio observations of XTE J1901+014 are strongly encouraged.

XTE J1751–305 Z. Wang and D. Chakrabarty, MIT, report that the candidate infrared

counterpart that they proposed for XTE J1751–305 (IAUC 7874) is now likely excluded. Further analysis of their J-band images of the field from Apr. 9 UT have led to a revised astrometric solution (0′′.5 rms) that places the candidate outside the 0′′.6-radius Chandra x-ray error circle (IAUC 7876), lying 0′′.8 west and 0′′.4 north of its center. No stars are detected within the error circle (limiting mag J = 20.5). The two detected stars closest to the Chandra position are the former candidate (0′′.9 distant) and one (J = 18.0 ± 0.5; blended with a neighbor to the southwest having J = 13.5) that is 1′′.2 distant.

C. B. Markwardt, University of Maryland and Goddard Space Flight Center, reports that XTE J1751–305 (IAUC 7867) has become undetectable by the RXTE PCA instrument. Since discovery, RXTE monitoring has shown a decrease in the x-ray flux, and as of Apr. 17.5 UT the 2–10-keV upper limit was ≈ 1 mCrab (due primarily to other nearby sources).

2002 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green